OpenStudy (anonymous):

So An inheritance of $31,000 will be divided sisters Hannah and Taylor. If Hannah is to receive $9000 less than Taylor, how much does both sisters receive?

OpenStudy (notshh):

what's up with that tho? totally unfair. hannah was prob adopted.

zepdrix (zepdrix):

lol poor Hannah :) So Taylor receives some amount, let's call that amount T. Then Hannah's amount, let's call that H, is 9000 less than that, or in other words, the amount T minus 9000. H=T-9000 And we also know that the total amount given to the sisters is 31000. So the amount given to Hannah `plus` the amount given to Taylor is 31000. H+T=31000

zepdrix (zepdrix):

We have a system of two equations, involving two unknowns.

zepdrix (zepdrix):

Substitution would be a good approach since we have H alone on one side of our first equation. \(\large\rm \color{orangered}{H=T-9000}\) We want to use this orange equation to replace the H in our second equation, \(\large\rm \color{orangered}{H}+T=31000\) Which becomes \(\large\rm \color{orangered}{T-9000}+T=31000\)

zepdrix (zepdrix):

Understand that step Kendy? What do you think? :)

OpenStudy (anonymous):

I don't know it wants to know how much both sister receive each?

OpenStudy (09395548052cs):

Let us analyze the second statement in the problem, carefully.. If Hannah is to receive $9000 less than Taylor... in other words, Tylor will receive 9000 dollars more than Hannah. KendyM25, my question is who has the less amount to recieve from the inheritance which is 31000? Who is probably the adopted one, is it Hannah or Tylor(according to notshh)? After that make a let statement just like, let x the inheritance of Hannah and let x+9000 the inheritance of Tyloy. Then make an equation to get the value of x.

OpenStudy (09395548052cs):

Hannah's inheritance: x Tylor's inheritance: x+9000 Total inheritance: 31000 Hannah's inheritace plus Tylor's inheritance= Total inheritance Equation: (x+9000) + x = 31000 Find the value of x and then substitute to get the amount, each will receive.